An ion accelerated through a potential difference of 115V experiences 

an increase in potential energy of 7.37×10-17J. 

Calculate the charge on the ion.


1
Expert's answer
2021-06-15T11:03:26-0400
"\Delta V=\dfrac{PE}{q},""q=\dfrac{PE}{\Delta V}=\dfrac{7.37\cdot10^{-17}\ J}{115\ V}=6.41\cdot10^{-19}\ C."